OpenStudy (anonymous):

The length of a rectangle is 5yrds more than twice its width, and the AREA of the rectangle is 88 . Find the dimensions of the rectangle

OpenStudy (anonymous):

Let \(l\) and \(w\) be the length and the width of the rectangle respectively, then \(l=2w+5\) and \((2w+5)w=88 \implies 2w^2+5w-88=0 \implies (2w-11)(w+8)=0\). Thus \(w=11/2\) and \(l=2(11/2)+5=16\).